OpenStudy (anonymous):

10. Challenge: Graph the equation with a diameter that has endpoints at (2, 2) and (-4, -2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

same idea as last time

jimthompson5910 (jim_thompson5910):

find the midpoint to get the center then find the distance from center to either endpoint to get the radius

OpenStudy (anonymous):

can you set it up :) its my last question n i'll finally be done :D

jimthompson5910 (jim_thompson5910):

you add up the corresponding coordinates, then divide by 2 to get the midpoint

jimthompson5910 (jim_thompson5910):

what do you get for the midpoint

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

ok what do you get when you add up the x coordinates

OpenStudy (anonymous):

can you set it for me? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

2 + (-4) = ??

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

divide that in half

OpenStudy (anonymous):

2/2=1

jimthompson5910 (jim_thompson5910):

-2/2 = -1

jimthompson5910 (jim_thompson5910):

so the x coordinate of the center is -1

OpenStudy (anonymous):

now what about y?

jimthompson5910 (jim_thompson5910):

add up the y coordinates

OpenStudy (anonymous):

2+-2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that adds to what

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

then you divide that in half to get 0/2 = 0

jimthompson5910 (jim_thompson5910):

So the center is (-1,0)

jimthompson5910 (jim_thompson5910):

now you use this center, and another point, to find the radius

OpenStudy (anonymous):

@jim_thompson5910 so whats my radius?

jimthompson5910 (jim_thompson5910):

use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) to find the distance from (-1,0) to (2, 2)

OpenStudy (anonymous):

d-2--1^2 2-0^2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((2-(-1))^2+( 2-0)^2) d = sqrt((2+1)^2+( 2-0)^2) d = sqrt((3)^2+(2)^2) d = ???

OpenStudy (anonymous):

9 +4 13

jimthompson5910 (jim_thompson5910):

so the radius is \[\Large \sqrt{13}\] units

OpenStudy (anonymous):

thts it?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you for all your help!! :D

jimthompson5910 (jim_thompson5910):

yw